# Dirac - von Neumann axioms in the setting of Continuous Model Theory

### v1

@@ Abstract
We recast the well-known axiom system of quantum mechanics used by physicists (the Dirac calculus) in the language of Continuous Logic. For the basic version of the axiomatic system we prove that along with the canonical continuous model the axioms have approximate finite models of large sizes, in fact the continuous model is isomorphic to an ultraproduct of finite models. We analyse the continuous logic quantifier corresponding to Dirac integration and show that in finite context it has two versions, local and global, which coinside on Gaussian wave-functions.
@@ 1 Introduction
The axiomatic formulation of quantum mechanics was introduced by Paul Dirac in 1930 [1] through a description of Hilbert space, and later developed with greater mathematical rigor in a monograph of 1932 by John von Neumann. Since 1930, Dirac went through several rewritings and new editions to refine his calculus to a level he considered satisfactory. In fact, von Neumann himself expressed his dissatisfaction not long after publication of his book, and spent considerable effort looking for alternatives [20]. Modern books present Dirac’s axioms in a succinct form, often omitting much of the technical detail.
@@
In section 2 we survey the axioms of quantum mechanics following [2]. Readers with a background in logic will notice that what physicists refer to as axioms is very far from what is a conventional set of axioms in a formal language even in its early form as presented e.g. by Hilbert’s axiomatisation of geometry [3]. We noted earlier, in [5] and in [7], that the language that Dirac introduced is that of continuous logic, CL, and the (rigged) Hilbert space which the axioms describe is a close analogue of cylindric algebra of Tarski, see [6], in a CL-version. In section 4 we go further and prove our main Theorem 4.7 stating that the continuous theory of Dirac - von Neumann given in Hilbert-space form (with integral operators) has two kinds of models: the canonical model $\mathbb{U}(\infty)$ based on wave-functions/continuous predicates on ${\mathbb{R}}^{m}$, and an asymptotic class of finite approximate models which we call H-structures, $\mathbb{U}(n).$ More precisely we prove that the canonical continuous model $\mathbb{U}(\infty)$ is isomorphic to a CL-ultraproduct of the finite $\mathbb{U}(n),$ $n\in{\mathbb{N}}.$
@@
As a matter of fact our $\mathbb{U}(n)$ are specific lattice models. Of course, a variety of lattice models have been in use in physics. They play an important role as toy models, as vehicles for specific calculations as well as modelling specific phenomena. However, the fact that the specific cyclic lattice models $\mathbb{U}(n)$ hosting the full set of integral operators represent in mathematically exact way Dirac’s axioms of quantum mechanics is novel. In particular, the Dirac integral is shown to correspond to a summation formula which is “local” in a certain well-defined sense, namely the summation domain has to be much shorter then the length of the full cycle. All unitary operators of the form $\mathrm{e}^{iL}$ where $L$ is self-adjoint in a form of a polynomial of $\mathrm{P}$ and $\mathrm{Q}$ with rational coefficients, are represented in the $\mathbb{U}(n).$
@@ 3 The axioms in the setting of Continuous Logic
Note that unitary operators of the form $\mathrm{e}^{iL},$ where $L$ is self-adjoint can mostly be represented in the integral form with respective kernels $\alpha.$ In particular, it is true when $L$ is a polynomial of the basic operators $\mathrm{P}$ and $\mathrm{Q},$ see [19].
@@ 4 Hilbert space formalism and H-structures
Let $\mathcal{H}$ be a rigged Hilbert space, $\mathcal{H}_{n}=\mathcal{H}^{\otimes n}$ and $\mathrm{O}$ a collection of linear operators on the $\mathcal{H}_{n}.$ Let $\mathcal{H}_{n}^{\mathrm{Def}}$ is a dense subspace of $\mathcal{H}_{n}$ closed under operators from $\mathrm{O}.$
We will often write $\mathcal{H}$ for the union of the tower $\mathcal{H}_{1}\subset\mathcal{H}_{2}\subset\ldots$ and similarly with $\mathcal{H}^{\mathrm{Def}}.$
Definition. An H-structure $(\mathbb{U},\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ is given by
- a universe $\mathbb{U},$ a complete metric space with a measure $\mu$ and metric $\mathrm{dist}(u_{1},u_{2}).$
In case $\mathbb{U}=\mathbb{U}(\mathfrak{n})$ is finite of size $\mathfrak{n},$ we identify
@@
with additive structure isomorphic to ${\mathbb{Z}}/\mathfrak{n}{\mathbb{Z}}$ and $\mathrm{dist}(u_{1},u_{2})=\sqrt{\frac{1}{\mathfrak{n}}}\cdot|u_{1}-u_{2}|$, the measure of a point is $\frac{1}{\sqrt{\mathfrak{n}}}.$
In case $\mathbb{U}=\mathbb{U}(\infty)$ is infinite, $\mathbb{U}(\infty):={\mathbb{R}},$ $\mathrm{dist}(u_{1},u_{2})=|u_{1}-u_{2}|$ and the measure is the Dirac’s delta-measure as determined in 3.5.
- collections $H_{n},$ $n\in{\mathbb{N}},$ of predicates
@@
each with a name $\psi$ from $\mathcal{H}_{n}^{\mathrm{Def}},$ continuous maps; $H_{n}$ has a structure of a $\mathbb{C}$}-linear space;
- an Hermitian inner product $\langle\psi_{1}|\psi_{2}\rangle;\ H_{n}\times H_{n}\to{\mathbb{C}}$ is defined for all $n$;
- quantifier
$\mathrm{E}:H_{n+1}\to H_{n},$
for all $n\in{\mathbb{N}}$, which is given as an integral operator
$\mathrm{E}:\psi\mapsto\int_{\mathbb{U}}\psi\,d\mu$
on the rigged Hilbert space $H_{n+1}.$
- a collection of linear operators
$L:H_{n}\to H_{m}$
named by symbols $L\in\mathrm{O}$:
- Weyl assumption: $\mathrm{O}$ contains a pair of (Weyl) operators $\mathrm{U}$ and $\mathrm{V},$ acting on $H_{1}$ and satisfying the commutation relation
$\mathrm{U}\mathrm{V}=q\mathrm{V}\mathrm{U}$
@@
$q=\left\{\begin{array}[]{ll}\mathrm{e}^{\frac{2\pi i}{\mathcal{N}}}\mbox{ if $\mathbb{U}$ is finite }|\mathbb{U}|=\mathcal{N}\\ \mathrm{e}^{2\pi i\mathfrak{h}}\mbox{ some }\mathfrak{h}\in{\mathbb{R}}_{>0}\end{array}\right$.
There is a canonical set of eigenvectors for $\mathrm{U}$
$\mathrm{Eig}\,\mathrm{U}=\{{\bf u}[r]\in H_{1}:\mathrm{U}{\bf u}[r]=q^{r}\cdot{\bf u}[r],\ r\in\mathbb{U}\}\ \ \ \langle{\bf u}[r]|{\bf u}[s]\rangle=\delta_{r,s}$
which form an orthonormal basis of $H_{1}.$ The action of $\mathrm{V}$ on the basis is defined as
[…]
There is a dual canonical set of eigenvectors for $\mathrm{V}$ with eigenvalues in $\mathbb{S}(\mathbb{U})\subset\mathbb{S}$
$\mathrm{Eig}\,\mathrm{V}=\{{\bf v}[p]\in H_{1}:\mathrm{V}{\bf v}[p]=q^{p}\cdot{\bf v}[p],\ p\in\mathbb{U}\}\ \ \ \langle{\bf v}[p]|{\bf u}[q]\rangle=\delta_{p,q}$
We call an infinite H-structure a continuous H-structure. Recall that $\mathbb{U}=\mathbb{U}(\infty)={\mathbb{R}}$ in this case.
4.3
There is a Fourier duality between the bases $\mathrm{Eig}\,\mathrm{U}$ and $\mathrm{Eig}\,\mathrm{V},$ in finite cases:
${\bf v}[p]=\frac{1}{\sqrt{\mathfrak{n}}}\sum_{r\in\mathbb{U}}q^{rp}{\bf u}[r];\ \ {\bf u}[r]=\frac{1}{\sqrt{\mathfrak{n}}}\sum_{p\in\mathbb{U}}q^{-rp}{\bf v}[p]$ (16)
or, in continuous/rigged Hilbert spaces form, in agreement with (5),
${\bf v}[p]=\frac{1}{\sqrt{2\pi}}\mathrm{e}^{-irp};\ \ {\bf u}[r]=\delta(r)=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}{\bf v}[p]\mathrm{e}^{ipr}dp$
In the setting of continuous model theory we represent the one-point characteritic functions ${\bf u}[r]$ as the limit of continuous bump-functions.
@@
We also consider asymptotic classes of H-structures with given signature $\mathcal{H}^{\mathrm{Def}},\mathrm{O},$ which are classes of H-structures with finite universe $\mathbb{U}(n),$ $n$ running in a subset $P\subseteq{\mathbb{N}}.$ (Similar to the class considered by Hrushovski in [15]). Then, given an ultrafilter $\mathcal{D}$ on $P$ and the first-order ultraproduct
$\mathbb{U}(\mathfrak{n})=\prod_{n\in\mathcal{D}}\mathbb{U}(n)$
we obtain a pseudo-finite version of H-structure. We study these below along with CL-ultraproducts.
We say that $(\mathbb{U},\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ represents the (possibly uncomplete) Hilbert space $(\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ if
$(H,\mathrm{O})\cong(\mathcal{H}^{\mathrm{Def}},\mathrm{O})$
via the naming correspondence.
@@
Proposition. Given a rigged Hilbert space $\mathcal{H}$ with an inner product subspace $\mathcal{H}^{\mathrm{Def}}$ which is dense in $\mathcal{H}$ and closed under the quantifier and a family of unitary operators $\mathrm{O}$ satisfying the Weyl assumtion, there is a unique H-structure $(\mathbb{U},\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ representing $(\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ .
Proof. Set $\mathbb{U}:=\mathrm{Eig}\,\mathrm{U}$ and define, for each $\psi\in\mathcal{H}^{\mathrm{Def}}_{n},$ an $n$-ary predicate $\psi$ to be the unique map $\psi:\mathbb{U}^{n}\to{\mathbb{C}}$ such that
$\psi(\bar{r})=\langle{\bf u}[\bar{r}]\,|\,\psi\rangle$
This is well-defined according to 3.6 (note that ${\bf u}[\bar{r}]=|\bar{r}\rangle,$ a position state in $\mathcal{H}_{n}$). $\Box$

### v2

@@ Abstract
We recast the well-known axiom system of quantum mechanics used by physicists (the Dirac calculus) in the language of Continuous Logic. For the basic version of the axiomatic system we prove that along with the canonical continuous model the axioms have approximate finite models of large sizes, in fact the continuous model is isomorphic to an ultraproduct of finite models. We analyse the continuous logic quantifier corresponding to Dirac integration and show that in finite context it has two versions, local and global, which coincide on Gaussian wave-functions.
@@ 1 Introduction
The axiomatic formulation of quantum mechanics was introduced by Paul Dirac in 1930 [1] through a description of Hilbert space, and later developed with greater mathematical rigor in a monograph of 1932 by John von Neumann. Since 1930, Dirac went through several rewritings and new editions to refine his calculus to a level he considered satisfactory. In fact, von Neumann himself expressed his dissatisfaction not long after publication of his book, and spent considerable effort looking for alternatives [20].
This discussion continues presently, see e.g. [12] questioning physicality of the complete Hilbert space $\mathcal{H}$ of quantum mechanics, effectively arguing for a dense subspace $\mathcal{H}^{\mathrm{Def}}$ of physically meaningfull states instead.
1.2
@@
In section 2 we survey the axioms of quantum mechanics following [2]. Readers with a background in logic will notice that what physicists refer to as axioms is very far from what is a conventional set of axioms in a formal language even in its early form as presented e.g. by Hilbert’s axiomatisation of geometry [3]. We noted earlier, in [5] and in [7], that the language that Dirac introduced is that of continuous logic, CL, and the (rigged) Hilbert space which the axioms describe is a close analogue of cylindric algebra of Tarski, see [6], in a CL-version. In section 4 we go further and prove our main Theorem 4.8 stating that the continuous theory of Dirac - von Neumann given in Hilbert-space form (with integral operators) has two kinds of models for the rigged Hilbert space $L^{2}({\mathbb{R}})$ which we call $\mathcal{H}$-structures:
- the canonical model $\mathbb{U}(\infty)$ based on wave-functions/continuous predicates on ${\mathbb{R}}^{m},$ and
- an asymptotic class of finite approximate models $\mathbb{U}(n).$
More precisely we prove that the canonical continuous model $\mathbb{U}(\infty)$ is isomorphic to a CL-ultraproduct of the finite $\mathbb{U}(n),$ $n\in{\mathbb{N}}.$
1.3
It is crucial that to represent elements of the rigged Hilbert space which are outside $L^{2}({\mathbb{R}}),$ including all-important position and momentum eigenvectors, one has to use the model-theoretic techniques of imaginary elements and imaginary sorts. Moreover, this technique allows to immediately generalise the setting of $L^{2}({\mathbb{R}}^{m})$, of quantum mechanics over the space ${\mathbb{R}}^{m}$, to the general $L^{2}({\mathbb{M}}),$ of quantum mechanics over a manifold $\mathbb{M}$} which is interpretable in $\mathbb{R}$}.
Note the key difference between the Hilbert space - reprepresntation of quantum mechanics and the CL-models representation: the latter has an explicit geoemtric universe $\mathbb{M}$} similar to the universe of Newtonian physics.
1.4
@@
As a matter of fact our $\mathbb{U}(n)$ are specific lattice models. Of course, a variety of lattice models have been in use in physics. They play an important role as toy models, as vehicles for specific calculations as well as modelling specific phenomena. However, the fact that the specific cyclic lattice models $\mathbb{U}(n)$ hosting the full set of integral operators represent in a mathematically exact way Dirac’s axioms of quantum mechanics is novel. In particular, the Dirac integral is shown to correspond to a summation formula which is “local” in a certain well-defined sense, namely the summation domain has to be much shorter then the length of the full cycle. All unitary operators of the form $\mathrm{e}^{iL}$ where $L$ is self-adjoint in a form of a polynomial of $\mathrm{P}$ and $\mathrm{Q}$ with rational coefficients, are represented in the $\mathbb{U}(n).$
@@ 4 Hilbert space formalism and H-structures
$\mathcal{H}$-structure.
Let $\mathcal{H}$ be a rigged Hilbert space, $\mathcal{H}_{n}=\mathcal{H}^{\otimes n}$ and $\mathrm{O}$ a collection of linear operators on the $\mathcal{H}_{n}.$ We will often write $\mathcal{H}$ for the union of the tower $\mathcal{H}_{1}\subset\mathcal{H}_{2}\subset\ldots$
Let $\mathcal{H}_{n}^{\mathrm{Def}}$ be subspaces of the $\mathcal{H}_{n}$ closed under operators from $\mathrm{O}.$
Given $(\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ we will associate with it a structure $(\mathbb{U};\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ in continuous model theory (sometimes called an $\mathcal{H}$-structure). $\mathbb{U}$ is the metric universe represented as a union of finite diameter domains, $(\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ is the signature.
There is a quantifier $\mathrm{E}_{x}$ on $(\mathbb{U};\mathcal{H}^{\mathrm{Def}},\mathrm{O})$, realised as the integral over a certain measure. Using the integral we define a Hermitian inner product on the linear spaces $\mathcal{H}^{\mathrm{Def}}_{n}.$ This gives us a structure of continuous model theory.
Then we can use the construction of imaginary elements and imaginary sorts to recover in $\mathbb{U}$ elements of spaces dual to $\mathcal{H}^{\mathrm{Def}}_{n}.$
We aim to represent in $(\mathbb{U};\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ enough of physically meaningful states, at least all-important position $|x\rangle$ and momentum $|p\rangle$ eigenvectors, and $\mathrm{O}$ to contain at least the unitary operators like the Weyl operators $\mathrm{e}^{i\mathrm{Q}}$ and $\mathrm{e}^{i\mathrm{P}}$ and general integral kernel operators.
4.3
Constructing an $\mathcal{H}$structure $\mathbb{U}$. The universe of $\mathbb{U},$ which we also call $\mathbb{U},$ is a complete metric space with a distinguished point $0,$ a measure $\mu$ and metric $\mathrm{dist}(u_{1},u_{2}).$
There are two cases of how the universe can be presented. In case $\mathbb{U}=\mathbb{U}(\infty)$ is infinite, define $\mathbb{U}(\infty):={\mathbb{R}},$ $\mathrm{dist}(u_{1},u_{2})=|u_{1}-u_{2}|$ and the measure is the Dirac’s delta-measure as determined in 3.5.
In case $\mathbb{U}=\mathbb{U}(\mathfrak{n})$ is finite of size $\mathfrak{n}$ (may assume $\mathfrak{n}$ even and $\sqrt{\mathfrak{n}}$ is an integer) we identify
@@
with metric given by $\mathrm{dist}(u_{1},u_{2})=\sqrt{\frac{2\pi}{\mathfrak{n}}}\cdot|u_{1}-u_{2}|$, the measure of a point $dx=\sqrt{\frac{2\pi}{\mathfrak{n}}},$ where $|u_{1}-u_{2}|$ is the natural integer valued distance in the cyclic group ${\mathbb{Z}}/\mathfrak{n}{\mathbb{Z}}.$
The universe $\mathbb{U}$ contains family of closed subdomains $[k,m]$ of finite diameters, $\in{\mathbb{R}}, k<mk<m$
$k,m]=\{u\in\mathbb{U}:\mathrm{dist}(k,u)+\mathrm{dist}(u,m)=m-k\}.$
and we assume that the end-points of the intervals are named.
In $\mathbb{U}(\infty)={\mathbb{R}}$ we consider Lipschitz-continuous $n$-ary predicates
$\psi:[k_{1},m_{1}]\times\ldots\times[k_{n},m_{n}]\to{\mathbb{C}}.$
The names of all such predicates is $\mathcal{H}^{\mathrm{Def}}.$ There is a natural uniform continuity modulus for each name $\psi.$
Equivalently, we can associate with each such name the map $\psi^{\prime}:\mathbb{U}^{n}\to{\mathbb{C}}$ which coincides with $\psi$ on $k_{1},m_{1}]\times\ldots\times[k_{n},m_{n}]$ and is 0 outside. Call such $\psi^{\prime}$ test functions with name $\psi.$
For each $n,$ let $\Phi_{n}$ be the $\mathbb{C}$}-linear space of functions $\phi:\mathbb{U}^{n}\to{\mathbb{C}}$ which are representable as finite linear combinations of test functions.
In case of the discrete $\mathbb{U}(\mathfrak{n})$ there is a natural embedding
$k,m]^{\mathbb{U}(\mathfrak{n})}\subset[k,m]^{\mathbb{U}(\infty)};\ u\mapsto\sqrt{\frac{2\pi}{\mathfrak{n}}}\,u$ (16)
and thus we define the predicate with the name $\psi$ on $\mathbb{U}(\mathfrak{n})$ as the restriction of $\psi$ to $k,m]^{\mathbb{U}(\mathfrak{n})}.$ This is obviously extendable to $n$-ary predicates.
We may use notation $\Phi_{n}(\infty),$ $\Phi_{n}(\mathfrak{n}),$ for spaces of predicates on $\mathbb{U}(\infty)$ and $\mathbb{U}(\mathfrak{n})$ respectively.
It is well known, for $\mathbb{U}(\infty)={\mathbb{R}},$ that $\Phi_{n}$ is a dense inner product subspace of $L^{2}({\mathbb{R}}^{n})$ which in many cases is taken to be $\mathcal{H}$ of the axioms. Moreover,
$\Phi_{n}\subset\mathcal{H}\subset\Phi^{*}_{n}$
where $\Phi^{*}_{n}$ is the dual to $\Phi_{n},$ that is the inner product space of anti-linear functionals on $\Phi_{n}.$
We introduce quantifier $\mathrm{E}:\Phi_{n+1}\to\Phi_{n}$ uniformly on all domains $m,m]^{n+1}$
$\mathrm{E}_{x}:\psi(x,y_{1},...,y_{n})\mapsto\int_{\mathbb{R}}\psi(x,y_{1},...,y_{n})dx=\int_{-m}^{m}\psi(x,y_{1},...,y_{n})dx$
In case of discrete $\mathbb{U}(\mathfrak{n})$ the quantifier is expressible by the summation formula
$\mathrm{E}_{x}:\psi(x,y_{1},...,y_{n})\mapsto\sum_{-m\leq k\sqrt{\frac{2\pi}{\mathfrak{n}}}<m}\psi(k\sqrt{\frac{2\pi}{\mathfrak{n}}},y_{1},...,y_{n})\cdot\sqrt{\frac{2\pi}{\mathfrak{n}}}$
where $k$ runs in ${\mathbb{Z}}\cap[-\frac{\mathfrak{n}}{2},\frac{\mathfrak{n}}{2}).$
The quantifier allows the definition of the inner product on $\Phi$ (with the natural extension to $\Phi_{n}$)
$\langle\psi|\phi\rangle:=\mathrm{E}_{x}\,\psi(x)\cdot\phi(x)^{*}$ (17)
where $\phi(x)^{*}$ stands for complex conjugation of $\phi(x).$
This implies that $\psi$ can be identified with the anti-linear functional
$\phi\mapsto\langle\psi|\phi\rangle,$
an element of $\Phi^{*}.$
In a slight extension of the standard continuous model theory setting we consider the continuous $n$-ary maps
@@
such that
$\psi=\bigcup_{m\in{\mathbb{N}}}\psi_{|[-m,m]^{n}}$ (18)
the union of basic continuous predicates.
We call these generalised predicates and consider them definable. This passes the main test of definablity. It follows from the representation (18):
Given a generalised predicate $\psi$ on $\mathbb{U}$ and an ultrafilter $\mathcal{D},$ the ultraproduct $\psi^{\mathcal{D}}$ of $\psi$ in the ultrapower $\mathbb{U}^{\mathcal{D}},$ is a generalised predicate.
On the other hand, for a generalised $\psi$ the inner product $\langle\phi|\psi\rangle$ with $\phi\in\Phi$ is well-defined since the effective domain of integration is bounded. So $\psi$ can be identified with the linear functional on $\Phi$
$\phi\mapsto\langle\phi|\psi\rangle,$
that is an imaginary element $\langle\psi|\in\Phi^{*}.$
We define momentum eigenfunctions ${\bf v}[p],$ $p\in{\mathbb{R}}$ as generalised continuous predicates on $\mathbb{U}(\infty)={\mathbb{R}}$
${\bf v}[p]:x\mapsto\frac{1}{\sqrt{2\pi}}\mathrm{e}^{\frac{ixp}{\mathbf{h}}}.$
Another way of representing ${\bf v}[p]$ is by considering the predicate ${\bf v}{\bf u}$ on $\mathbb{U}^{2}(\infty):$
${\bf v}{\bf u}:{\mathbb{R}}^{2}\to{\mathbb{C}};\ \ (x,p)\mapsto\langle x|p\rangle=\frac{1}{\sqrt{2\pi}}\mathrm{e}^{-\frac{ixp}{\mathbf{h}}}$
This defines the family
$\mathbb{V}=\{{\bf v}[p]:p\in\mathbb{U}\}$
as an imaginary sort (in the sense slightly broader than that in [14] ) with a metric given by
$\mathrm{dist}({\bf v}[p_{1}],{\bf v}[p_{2}])=\mathrm{dist}(p_{1},p_{2}).$
It is clear that $\mathbb{V}\cong\mathbb{U}$ as metric spaces. $\mathbb{V}$ is an imaginary sort in $(\mathbb{U};\mathcal{H}^{\mathrm{Def}}).$
Define, for $x\in\mathbb{U},$ the maps
${\bf u}[x]:\mathbb{V}\to{\mathbb{C}};\ {\bf v}[p]\mapsto{\bf v}[p](x)=\frac{1}{\sqrt{2\pi}}\mathrm{e}^{\frac{ixp}{\mathbf{h}}}$
These are generalised continuous predicates on sort $\mathbb{V}.$ Moreover, the above can be extended to a unique linear functional on $\Phi$ and so ${\bf u}[x]\in\Phi^{*}.$
${\bf u}[x]:\Phi\to\Phi;\ \psi\mapsto\psi(x)=\langle\psi|x\rangle$ (19)
This determines ${\bf u}[x]$ (or $|x\rangle$ in bra-ket notations) as imaginary elements and
$\mathbb{U}^{\prime}:=\{{\bf u}[x]:x\in\mathbb{U}\}$
a sort of imaginaries in a natural isomorphism with $\mathbb{U}.$ The ${\bf u}[x]$ are not in $\mathcal{H}$ but in Dirac calculus one defines their norm to be the symbolic Dirac $\delta(0).$ In the linear space $\Phi^{*}$ one can equivalently introduce the inner product
$\langle{\bf u}[x]|{\bf u}[y]\rangle:=\delta_{x,y}\ (\mbox{Kronecker }\delta).$
In this sense $\mathbb{U}^{\prime}$ is an orthogonal basis for $\Phi$ in the sense that, for each $\psi\in\Phi,$
$\psi=\mathrm{E}_{x}\psi(x){\bf u}[x]$ (20)
since, by linearity of $\mathrm{E},$
$\langle\,\mathrm{E}_{x}\psi(x){\bf u}[x]\,|\,{\bf u}[y]\,\rangle=\psi(y).$
Operators. There are two definable ways of introducing operators on $\Phi.$
1.Integral kernel operators. Let $\kappa:\mathbb{U}^{2}\to{\mathbb{C}}$ be a generalised predicate in variable z,x.{z},x. Then for all $\psi\in\Phi_{n}$ in variables $x,\bar{y}$
$L_{\kappa}:\psi\mapsto\mathrm{E}_{x}\,\kappa({z},x)\cdot\psi(x,\bar{y})$
is well-defined and determines a linear operator $\Phi_{n}\to\Phi_{n}.$
2.Operators defined by their actions on bases.
We assume that $\mathrm{O}$ contains a pair of (Weyl) operators $\mathrm{U}$ and $\mathrm{V}$ on $\Phi$ defined on the ${\bf u}[r],\ r\in\mathbb{U}:$
$\mathrm{U}:{\bf u}[r]\mapsto\mathrm{e}^{ir\nu}\cdot{\bf u}[r]\mbox{ and }\mathrm{V}:{\bf u}[r]\mapsto{\bf u}[r+\mathbf{h}],\mbox{ if }\mathbb{U}=\mathbb{U}(\infty)\\$ (21)
@@
$\nu=1$ if $\mathbb{U}=\mathbb{U}(\infty)$, and $\nu=\frac{2\pi}{\mathfrak{n}}$ if $\mathbb{U}=\mathbb{U}(\mathfrak{n}).$
$\mathbf{h}=\hbar$ a positive real, if $\mathbb{U}=\mathbb{U}(\infty)$, and $\mathbf{h}=h_{\mathfrak{n}},$ a positive integer, if $\mathbb{U}=\mathbb{U}(\mathfrak{n}).$
(Note the correspondence (16) between units of $\mathbb{U}(\infty)$ and $\mathbb{U}(\mathfrak{n})$. We will later assume that $\sqrt{\mathfrak{n}}$ is an integer and $\lim_{\mathfrak{n}\to\infty}\frac{h_{\mathfrak{n}}}{\sqrt{\mathfrak{n}}}=\hbar$).
The following commutation relation then holds:
$\mathrm{U}\mathrm{V}=\mathrm{e}^{i\nu\mathbf{h}}\mathrm{V}\mathrm{U}.$
Using (20) we then get
$\begin{array}[]{ll}\mathrm{U}\psi=\mathrm{E}_{x}\psi(x)\cdot\mathrm{e}^{i\nu x}{\bf u}[x]\\ \mathrm{V}\psi=\mathrm{E}_{x}\psi(x-\mathbf{h}){\bf u}[x]\end{array}$
Thus $\mathrm{U}$ transforms function $\psi(x)$ on $[k,m]$ to $\psi(x)\cdot\mathrm{e}^{i\nu x}$ and $\mathrm{V}$ shifts $\psi(x)$ on $[k,m]$ to $\psi(x-\mathbf{h})$ defined on $k+\mathbf{h},m+\mathbf{h}]$.
On the dual set of vectors ${\bf v}[p],\ p\in\mathbb{U}:$
$\mathrm{V}:{\bf v}[p]\mapsto\mathrm{e}^{-i\nu p\mathbf{h}}{\bf v}[p]\mbox{ and }\mathrm{U}:{\bf v}[p]\mapsto{\bf v}[p-1].$
@@
Remarks.
1. There is a Fourier duality operator between the bases ${\bf u}[r]$ and ${\bf v}[p]$ of $\Phi^{*}.$
In the finite case:
${\bf v}[p]=\frac{1}{\sqrt{\mathfrak{n}}}\sum_{r\in\mathbb{U}(\mathfrak{n})}\mathrm{e}^{\frac{2\pi ih_{\mathfrak{n}}rp}{\mathfrak{n}}}{\bf u}[r];\ \ {\bf u}[r]=\frac{1}{\sqrt{\mathfrak{n}}}\sum_{p\in\mathbb{U}(\mathfrak{n})}\mathrm{e}^{-\frac{2\pi ih_{\mathfrak{n}}pr}{\mathfrak{n}}}{\bf v}[p]$ (22)
and in the continuous case, in agreement with (5),
${\bf v}[p]=\frac{1}{\sqrt{2\pi}}\mathrm{e}^{irp};\ \ {\bf u}[r]=\delta(r)=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}{\bf v}[p]\mathrm{e}^{-ipr}dp$
2. Note that unitary operators of the form $\mathrm{e}^{iL},$ where $L$ is self-adjoint can be represented in the integral form with respective kernels $\kappa$ when $L$ is a polynomial of the basic operators $\mathrm{P}$ and $\mathrm{Q},$ see [19].
3. The time evolution operator $\mathrm{e}^{iHt}$ for the quantum harmonic oscillator of frequency $\omega$ can be represented by an integral operator with kernel
$\sqrt{\frac{\omega}{2\pi i\hbar\sin\omega t}}\exp i\omega\frac{(x_{0}^{2}+x^{2})\cos\omega t-x_{0}x}{2\hbar\sin\omega t}$
@@
Proposition. Given the $\mathcal{H}$-structure $(\mathbb{U}(\infty);\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ constructed in 4.3, there is a unique rigged Hilbert space $\mathcal{H}$ which contains a dense inner product subspace of continuous predicates on bounded domains isomorphic to $\mathcal{H}^{\mathrm{Def}}$ with operators $\mathrm{O}.$
Given the $\mathcal{H}$-structure $(\mathbb{U}(\mathfrak{n});\mathcal{H}^{\mathrm{Def}},\mathrm{O})$ constructed in 4.3, the space of predicates on bounded domains with respect to the inner product is an $\mathfrak{n}$-dimensional Hilbert space with operators $\mathrm{O}.$
Proof. This is the content of 4.3.